OpenStudy (anonymous):

What is the x-intercept of the graph: 2x + 4y = -10

Directrix (directrix):

The x-intercept of the line which is the graph of 2x + 4y = -10 is the point where the graph cuts the x-axis. All points on the x-axis have a y-coordinate of 0. So, take the equation 2x + 4y = -10 and substitute 0 for y. Then, solve the resulting equation for x. That is the x-intercept. @AllenH666 Post you answer and we will compare, if you like.

OpenStudy (anonymous):

so you would get 5 for the x coordiante

OpenStudy (anonymous):

-5 actually

OpenStudy (anonymous):

how so?

Directrix (directrix):

@lilcaboboy ---> how so? All points on the x-axis have a y-coordinate of 0. So, take the equation 2x + 4y = -10 and substitute 0 for y. Then, solve the resulting equation for x. That is the x-intercept. 2x + 4y = -10 2x + 4*0 = -10 2x + 0 = -10 2x = -10 (1/2) * (2x) = (1/2)*(-10) x = ?